### v2

# dfigure[$\NewEnvirondfigure*[1]$

Abstract
Ultracold fermionic atoms in optical lattices offer pristine realizations of Hubbard models [1], which are fundamental to modern condensed matter physics [2, 3]. Despite significant advancements [4, 5, 6], the accessible temperatures in these optical lattice material analogs are still too high to address many open problems [7, 8, 9, 10]. Here, we demonstrate a several-fold reduction in temperature [11, 6, 12, 13], bringing large-scale quantum simulations of the Hubbard model into an entirely new regime. This is accomplished by transforming a low entropy product state into strongly-correlated states of interest via dynamic control of the model parameters [14, 15], which is extremely challenging to simulate classically [10]. At half filling, the long-range antiferromagnetic order is close to saturated, leading to a temperature of $T/t=0.05_{-0.05}^{0.06}$ based on comparisons to numerically exact simulations. Doped away from half-filling, it is exceedingly challenging to realize systematically accurate and predictive numerical simulations [9]. Importantly, we are able to use quantum simulation to identify a new pathway for achieving similarly low temperatures with doping. This is confirmed by comparing short-range spin correlations to state-of-the-art, but approximate, constrained-path auxiliary field quantum Monte Carlo simulations [16, 17, 18]. Compared to the cuprates [19, 20, 2], the reported temperatures correspond to a reduction from far above to below room temperature, where physics such as the pseudogap and stripe phases may be expected [19, 21, 22, 23, 3, 24]. Our work opens the door to quantum simulations that solve open questions in material science, develop synergies with numerical methods and theoretical studies, and lead to discoveries of new physics [8, 10].
Introduction
The Hubbard model is a paradigmatic description of strongly-correlated electrons that is central to our understanding of quantum materials [3]. After decades of concentrated study, it is strongly believed that this model hosts a variety of exotic quantum states [25, 26], including the intriguing phases observed in the cuprate high-temperature superconductors [2, 22, 3]. Despite its apparent simplicity, solving the two-dimensional, square-lattice Hubbard model has proved an outstanding challenge. In recent years, building on decades of algorithmic development, combined use of the most advanced numerical methods has established certain aspects of the model’s properties [27, 21, 28, 29, 23, 30, 24]. Nevertheless, a full description of the low-temperature phase diagram remains beyond the reach of theoretical tools and computational methods [9].
Its broad applicability and numerical intractability make the Hubbard model an ideal candidate for quantum simulation [8, 10]. In particular, ultracold fermionic atoms in optical lattices can natively implement the requisite fermionic degrees of freedom, and serve as a pristine model system for strongly-correlated electrons moving and interacting in a crystalline lattice [1]. The Hubbard model is described by the geometry of the lattice, a tunneling amplitude $t$ between neighboring sites, and a contact interaction energy $U$, which can all be widely programmed in cold-atom systems. However, the enlarged length and time scales in cold-atom systems lead to very small energy scales, making it challenging to reach regimes corresponding to real materials at low temperatures [7]. The lowest reported experimental temperatures of $T=0.25(2)t$ [11, 6, 12, 13] (Fig. Acknowledgementsa) in two-dimensional Hubbard quantum simulators with intermediate interaction strengths of $U/t\simeq 8$ correspond to an effective temperature of approximately $700$ K in the cuprates by normalizing temperatures with superexchange energy $J=4t^{2}/U$ [19, 20, 2] (see Methods). One needs to reach well below room temperature in order to investigate properties relevant to cuprates, including pseudogap behavior, stripe order, and unconventional superconductivity [19, 21, 22, 23, 3, 24]. New cooling schemes are therefore needed for cold-atom-based platforms to offer competitive simulations of quantum materials.
tp
tp
tp
In this work, we demonstrate a substantial reduction in the temperatures achievable in cold-atom-based quantum simulations of the Hubbard model on a two-dimensional square lattice. At half-filling, we achieve temperatures of $T=0.05_{-0.05}^{0.06}\,t$ with $U/t\simeq 8$ in a large system comprised of $\simeq 340$ lattice sites, which corresponds to a several-fold improvement of the state of the art [11, 6, 12, 13]. We compare the measured spin correlations, which extend out to long range, to exact numerical results, and find excellent agreement.
The behavior of the Hubbard model is not fully understood at finite doping, which has hindered the development of cooling techniques in this regime, even at the level of theoretical proposals. Here, we experimentally propose and demonstrate a novel pathway to realize low temperatures with finite values of doping $\delta$ between $2\%$ and $21\%$. Detailed comparisons with state-of-the-art approximate numerical computations [16, 17, 18] indicate that our temperatures are $T\lesssim 0.1\,t$, comparable to those achieved at half-filling. These temperatures correspond to a reduction to significantly below room temperature in cuprates.
Our scheme hinges on the efficient preparation of product states with extremely low entropies  [31]. Seminal proposals suggested that these states can be adiabatically connected to strongly-correlated states of interest to reach low temperatures [32, 14, 15, 33, 34]. Despite experimental progress with bosonic systems [35, 36, 37], practical implementations of these schemes for fermions have been challenging beyond proof-of-principle realizations at half-filling [38, 31, 39, 40]. This holds especially true in large Hubbard systems, where finite-time, out-of-equilibrium dynamics are out of reach of numerical simulations [41, 42]. Our work therefore showcases how quantum simulators can experimentally address such optimization problems.
Experimental scheme
The spirit of our protocol is similar to most cooling cycles. In a first generalized compression step, entropy is extracted from the system of interest to a reservoir in thermal contact by reducing the density of states of the system [43, 14, 31]. After the system is isolated from the reservoir, increasing the density of states by generalized expansion decreases the temperature. In practice, our scheme involves initializing cold atoms in a low entropy band insulator (BI), and transforming the BI into a strongly-correlated state of interest (Fig. Acknowledgementsb) [31, 39].
The BI can be prepared with extremely high fidelity because the chemical potential lies within the band gap featuring low density of states, which allows for efficient entropy redistribution to a metallic reservoir [43, 14, 11, 31, 36]. We load a spin-balanced mixture of fermionic 6Li atoms in the lowest two hyperfine states into an optical lattice (long-spacing lattice). We set the magnetic bias field at 550550G close to a broad Feshbach resonance to minimize interaction strengths for BI formation. Programmable optical potentials created by two separate Digital Micromirror Devices (DMDs) first confine about $340$ atoms into a BI covering $170$ sites, then isolate the BI from the reservoir. We estimate a low initial entropy per particle of $s=0.025(4)k_{B}$ [31] based on a measured singly-occupied site density of $n_{s}=0.7(2)\%$ in a central disk of $\simeq 110$ sites, although an accurate estimation is complicated by gradients at the edge of the confining potential due to finite optical resolution (see Methods).
Keeping the subsequent transformation slow relative to the relevant many-body timescales is crucial to realizing low final temperatures. This is particularly challenging as this protocol involves a substantial change in lattice filling from two in the BI to around one atom per site in strongly-correlated Hubbard systems. Such a massive change in the state typically involves very slow many-body timescales, which requires excellent local control of density [38, 31, 36, 39, 40]. To address this challenge, we take advantage of an optical lattice whose geometry is dynamically adjustable [44, 45, 12]. In particular, we continuously split single lattice sites into double-wells, and subsequently connect these double-wells together into a square lattice, realizing a scheme similar to the one proposed in [15]. This doubles the number of sites at fixed total atom number (Fig. Acknowledgementsc,d and Extended Data Fig. Competing Interests), and naturally converts the BI into half-filled Fermi systems. However, this doubling is not exact in the experiment due to variations of the optical potentials. To precisely reach the target filling, our protocol further relies on tunable interactions and DMD potentials to ensure controlled preparation, isolation and expansion of the system of interest (see Methods).
Splitting an insulator
Dynamically changing the lattice geometry leads to rich physics even at very large interactions of $U/t=18.6(8)$, where the half-filled state resulting from splitting the BI is a Mott insulator with a large charge gap. Here, the Hubbard model simplifies to a spin model with antiferromagnetic Heisenberg coupling $J=4t^{2}/U$. In the thermodynamic limit, this Heisenberg model is known to host a quantum phase transition between a disordered phase in the dimerized lattice, described as a product of isolated singlet states within double wells, and a Néel ordered phase in the square lattice with long-range antiferromagnetic correlations (Fig. Acknowledgementsa) [46, 47]. In our experiment, which is a finite-sized system, the many-body spin gap is expected to decrease monotonically during splitting, with a minimal value of $\Delta_{H}\propto J/L^{2}$ corresponding to the Anderson tower of states with broken symmetry [48]. These two lattice geometries are continuously connected by tuning the coupling between double wells, which preserves total spin $S_{\mathrm{tot}}=0$ and enables efficient low-entropy state preparation [15]. We parametrize the transformation of lattice geometry by a single parameter $\alpha$ describing the ratios of tunneling amplitudes within and between double wells (see Fig. Acknowledgementsa and Methods).
To track how spin order changes with lattice geometry, we experimentally measure the spin correlation function in the $z$ direction between sites at positions $\mathbf{r}$ and $\mathbf{r+d}$ [12]:
$C^{zz}_{\mathbf{d}}(\mathbf{r})$ $=\frac{1}{S^{2}}\left(\langle S^{z}_{\mathbf{r}}S^{z}_{\mathbf{r+d}}\rangle-\langle S^{z}_{\mathbf{r}}\rangle\langle S^{z}_{\mathbf{r+d}}\rangle\right).$ (1)
In the $\alpha\rightarrow 0$ limit of disconnected dimers, which corresponds to the BI, we only detect saturated spin correlations between nearest-neighbor sites, $|\mathbf{d}|=1$ within the dimers and consistent with 0 everywhere else (Fig. Acknowledgementsb). As these dimers are connected, spin correlations start to grow between the dimers at $\alpha\simeq 0.7$ and become uniform and isotropic in the square lattice $\alpha=1.0$.
To probe spin correlations at longer bond distances $|\mathbf{d}|>1$, we obtain $C^{zz}_{\mathbf{d}}$ by averaging $C^{zz}_{\mathbf{d}}(\mathbf{r})$ over a Mott insulating region within a central disk of radius $r=6$, and account for spatial inversion symmetry (see Methods). We confirm that spin correlations are antiferromagnetic and localized on the intra-dimer bonds in the dimerized lattice $\alpha=0.3$, and that they extend to long range in the square lattice with $\alpha=1.0$ (Fig. Acknowledgementsc).
The nature of magnetic order during the lattice transformation is captured by the staggered magnetization $m^{z}$ [11]. We compute $(m^{z})^{2}$ by averaging the sign-corrected spin correlations $(-)^{|d_{x}+d_{y}|}C^{zz}_{\mathbf{d}}$ up to a cutoff distance $d_{\mathrm{max}}$ (see Methods, Eq. 3). We observe an increase of $(m^{z})^{2}$ around $\alpha\simeq 0.7$ (Fig. Acknowledgementse), consistent with the range of critical values reported in previous numerical studies of dimerized Heisenberg models [46, 47]. The experimental data is also consistent with quantum Monte Carlo simulations performed at a fixed temperature of $T=0.5J=0.12t$.
We experimentally check ramp adiabaticity and confirm that the spin correlations have saturated by varying the ramp duration, which could be limited by heating. Surprisingly, we find the strengths of the DMD potentials confining the Mott insulator, which may lead to atom transport into and out of the central region, can strongly affect the final spin correlations, and therefore temperatures (see Methods and Extended Data Fig. Competing Interests). These measurements suggest charge transport may be crucial to temperature optimization [49, 50] in addition to spin dynamics of the idealized model.
tp
Ultralow temperatures in the strongly-correlated regime
At intermediate interaction strengths of $U/t\simeq 8$ (Fig. Acknowledgementsa), which are most relevant to cuprate physics, the Hubbard model remains an antiferromagnet at half-filling, and can thus be reached by the same adiabatic ramp as in the Heisenberg regime. However, due to the reduced interaction strength, the system has increased compressibility, which can lead to enhanced charge transport. As a result, we find that the harmonic confinement produced by the lattice laser intensity profile results in density variation across the system after splitting. Thus, if the central density is to remain $n=1$, the edge of the system must be at $n<1$. Since the entirety of the system began as a BI, this means atoms at the edge of the trap must be spilled out as the insulating gap is reduced during splitting. We achieve this by dynamically adjusting the strength of the DMD potentials (Fig. Acknowledgementsb). During this step, we empirically find that it is helpful to allow for transport to occur at low interaction strengths, which could be related to the fact that a Fermi liquid is more compressible than a Mott insulator (Fig. Acknowledgementsa) [50, 41, 42]. The timing and shape of the above ramps are extensively optimized based on the final temperature.
With the optimized experimental sequence, we observe antiferromagnetic spin correlations $C^{zz}_{\mathbf{d}}$ extending across the entire half-filling region with a radius $r=5$ for an interaction strength of $U/t=8.3(2)$ (Fig. Acknowledgementsc, see Methods). We compare these experimentally measured correlations after sign-correction and azimuthal average $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ to numerical results from simulations using finite-temperature determinant quantum Monte Carlo (DQMC) and ground-state auxiliary-field quantum Monte Carlo (AFQMC) methods, under open boundary conditions. At half-filling, the numerical results are in principle exact, although care must be taken to obtain accurate and unbiased spin correlations, especially at larger bond distances $d$ (see Methods). We find the measured $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ are close to saturation and consistent with numerical data at $T/t\in[0,0.1]$ (Fig. Acknowledgementsd). These strong, long-ranged correlations translate to a narrow peak in the spin structure factor $S(\mathbf{q})$ at quasimomentum $\mathbf{q}=(\pi,\pi)$ (Fig. Acknowledgementse), which is obtained as the Fourier transformation of $C^{zz}_{\mathbf{d}}$ (see Methods). Comparing numerical DQMC data to the experimental value of the squared staggered magnetization $(m^{z})^{2}$ computed up to a cutoff $d_{\text{max}}=6$ allows us to extract a temperature of $T/t=0.05_{-0.05}^{0.06}$ (Fig. Acknowledgementsf and Extended Data Fig. Competing Interests).
Doping cold antiferromagnets
The most intriguing and poorly understood physics of the Hubbard model resides in the doped regime. At low temperatures of $T/t<0.2$, and with intermediate interactions of $U/t\simeq 8$, the doped Hubbard model is extremely challenging. Unlike for half-filling, no numerically exact results for spin correlations $C^{zz}_{d}$ are available in this regime for sufficiently large system sizes, which have systematically explored parameters such as temperature and interaction. Extending our cooling scheme to the above regime is therefore very valuable to advance our understanding, since quantum simulation can offer direct measurements, potentially adding a powerful new element in conjunction with approximate numerical simulations and analytic theory.
An outstanding challenge is that dopants must be coherently introduced and delocalized into the nearly half-filled system after splitting to avoid detrimental heating [49, 50, 31]. To address this challenge, we must adapt our cooling scheme to reach a final state with a central density $n<1$. We begin with a band insulator and then double the density of lattice sites, resulting in a central half-filled region. Additional expansion of atoms out of this region is then required to achieve $n<1$. In prior work, however, such expansion incurred substantial heating [31], resulting in final temperatures of $T/t>0.4$. In this work, we circumvent these challenges by performing the expansion in a shallow lattice, which increases the tunneling amplitudes and reduces interaction strengths, facilitating the efficient transport of particles. At the same time, this decreases dissipative effects from the lattice light, reducing the background heating rate.
Our cooling scheme for doped systems is thus as follows (Fig. Acknowledgementsa,b): after forming the BI, we reduce the lattice depth by $80\%$ and perform splitting. The large tunnelings and weak interactions transform the BI into a Fermi liquid. We then expand the atoms by reducing the strength of the confining DMD potential, while ramping the magnetic bias field to its final value of 590590G. The depth of the square lattice is then ramped up to its final value, reaching the strongly-correlated regime of $U/t=8.0(3)$ (see Methods). At the same time, the strengths of the DMD potentials are adjusted to minimize transport during lattice loading [49, 50] (see Methods).
We report in Fig. Acknowledgementsd the spin correlations $C^{zz}_{\mathbf{d}}$ averaged over a region-of-interest (ROI) of radius $r=3$. We observe antiferromagnetic spin correlations extending over this entire region for hole dopings of up to $\delta=10\%$. To understand how spin correlations evolve with increasing doping, we plot the azimuthal average of the sign-corrected spin correlations $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$. Larger doping induces stronger relative suppression of spin correlations at longer bond distances, which reduces the range of antiferromagnetism. Interestingly, the absolute magnitude of correlation reduction seems insensitive to bond distances, which suggests doping decreases a long-range offset instead of affecting the short-range structure (Fig. Acknowledgementse).
We compare the experimentally measured spin correlations systematically against numerical results obtained with the constrained-path (CP) auxiliary field quantum Monte Carlo (AFQMC) method [16, 17, 18], which is a state-of-the-art approximate numerical technique. The spin correlations predicted by CP-AFQMC grow monotonically as temperature decreases. For a hole doping range of $\delta\in[2\%,21\%]$, we find good agreement between the experimentally measured nearest-neighbor spin correlations $C^{zz}_{1}$ and CP-AFQMC simulations at temperatures of $T/t\in[0,0.1]$. Simulations at $T/t=0.15$ exhibit significantly weaker correlations (Fig. Acknowledgementsf).
tp
Interestingly, for bond distances of $d>1$, although the agreement is good on the scale of the variation between modestly low temperatures ($T/t=0.25$) and the ground state, the experiment shows stronger correlations than predicted by CP-AFQMC. It is not quite at the level of the agreement found at half-filling between experimental data and numerically exact results for all bond distances. We verify that our observations do not depend on the weak spatial variation of doping due to harmonic confinement by changing the radius $r$ of the experimental ROI, and further confirm that the system is at thermal equilibrium (see Extended Data Fig. Competing Interests). Uncertainty in the calibration of experimental parameters (see Methods) is insufficient to explain the deviation, but other details in Hamiltonian difference (e.g. higher order corrections) could potentially contribute, especially at lower temperatures. The discrepancy may also hint at an underestimation of long-range correlations by CP-AFQMC. This would be consistent with data at elevated temperatures of $T/t=0.25,0.33$, where numerically exact DQMC results seem to be more consistent with experimentally measured spin correlations, and CP-AFQMC results show slightly lower correlations (see Methods and Extended Data Fig. Competing Interests). These systematic comparisons are an example of the synergies that now exist between quantum simulations and numerics in the regime where obtaining exact numerical results are incredibly challenging.
Previous attempts [31] at adiabatic state engineering in the Hubbard model suffered from excessive heating during the transformation from BI to Hubbard systems. Our scheme departs from such attempts in two ways: the use of a geometry tunable optical lattice enables us to halve the filling of the system without macroscopic density redistribution, and working at weak interactions in a shallow optical lattice minimizes heating effects during the remaining expansion. To probe the the importance of the latter choice, we prepare a BI in a square lattice and expand it to a half-filled system in the same lattice by shaping the DMD potentials, and measure the temperature of the final state as a function of the expansion duration $\tau$ and the normalized lattice depth $\eta$ during expansion ($\eta=1$ indicates full depth).
To characterize the adiabaticity of the expansion, we scan the expansion duration at fixed depth $\eta=0.2$ (Fig. Acknowledgementsa). We observe a rapid increase in the temperature at short expansion durations, indicating $\tau=20\mathrm{ms}\simeq 35\hbar/t$ as a critical timescale below which diabatic heating limits the temperature. As the lattice depth is increased, the tunneling energy decreases, requiring longer expansion durations to remain adiabatic. To probe the effects of lattice depth, we therefore measure the final temperature as a function of lattice depth at a fixed normalized expansion duration $\tau\simeq 60\hbar/t$, in which $t$ is computed from $\eta$ via a bandstructure calculation (see Methods). As shown in Fig. Acknowledgementsb, the final temperature rapidly rises when the lattice depth is increased beyond a critical value $\eta\simeq 0.6$, which cannot be due purely to single-particle adiabaticity effects. Rather, we attribute it to a combination of interaction-induced heating during particle transport, which worsens as $U/t$ is increased, and dissipative effects induced by the lattice light and DMD light, which worsen due to the increase of $\eta$ and $\tau$.
The interaction-induced heating may be related to theoretical and experimental studies which have indicated that transport in strongly-correlated systems can experience an exponential slowdown [51, 52, 42] and induce heating from density redistribution [49, 50]. We note that the interaction strength $U/t\simeq 1.5$ at the critical lattice depth indicated in Fig. Acknowledgementsb is similar to the interaction at which a drastic increase in transport timescales was observed in [52]. These measurements highlight the importance of performing the expansion step in a shallow lattice, in addition to the use of a superlattice to double the density of sites.
Outlook
In this work, we have demonstrated a substantial reduction in the temperatures achievable in large-scale quantum simulations of the Hubbard model, both with and without doping. Through careful dynamical tuning of parameters in a programmable optical lattice, we are able to transform a trivial low entropy state into a cold, strongly-correlated quantum many-body state in equilibrium. Critically, the dynamics involved in this transformation are exceedingly difficult to tackle via analytical techniques, or to simulate classically, and so must be optimized empirically using the quantum simulator itself. The above approach is broadly applicable to situations where one can prepare low entropy product states with high fidelity, for example with optical tweezers [53]. These techniques can also be extended to other lattice geometries, including triangular and kagome lattices, which may host quantum spin liquids [26], and square lattices with diagonal tunneling $t^{\prime}$, which may host unconventional superconductivity [54].
The temperatures and system sizes achieved in this work likely allow one to enter phases of the Hubbard model that have not yet been explored in cold atom quantum simulators, including phases involving charge order [22, 21, 23, 3, 6]. Recent advancements involving comparisons of different approximate numerical simulations in equilibrium [27, 28] have shed some light on these behaviors, including the interplay between the stripe and pseudogap phases and $d$-wave superconductivity [24, 54]. However, a complete microscopic understanding has remained elusive. Quantum simulations in the appropriate regimes offer a unique opportunity to study features that are challenging to simulate classically, including spectral properties and real time dynamics. Additionally, in equilibrium, quantum simulations provide a valuable extra data point that complements approximate numerical simulations by taking dramatically different assumptions and approximations.
While advanced numerical simulations have long informed the design and operation of cold atom-based quantum simulators, this work opens the possibility of a fruitful exchange, wherein results obtained via quantum simulation can also be used to develop and benchmark more efficient and accurate numerical techniques. The benefits of combining classical and quantum tools go beyond simply serving as mutual benchmarks. The success of the preparation scheme in this work suggests that, in the future, hybrid classical-quantum algorithms that use results obtained from a quantum simulator to optimize state preparation [55, 56] could provide significant advantages, leading to lower temperatures, and possibly a definitive understanding of the Hubbard model in and out of equilibrium.
Acknowledgements
We thank Daniel Greif, Geoffrey Ji and Christie Chiu for early experimental contributions, Annabelle Bohrdt, Eugene Demler, Tilman Esslinger, Antoine Georges, Fabian Grusdt, Ehsan Khatami and Martin Zwierlein for insightful discussions, and Yuan-Yao He, Yiqi Yang and Zhou-Quan Wan for help with our calculations. We acknowledge support from the Gordon and Betty Moore Foundation, Grant No. GBMF-11521; National Science Foundation (NSF) Grants Nos. PHY-1734011, OAC-1934598 and OAC-2118310; ONR Grant No. N00014-18-1-2863; the Department of Energy, QSA Lawrence Berkeley Lab award No. DE-AC02-05CH11231; QuEra grant No. A44440; ARO/AFOSR/ONR DURIP Grants Nos. W911NF-20-1-0104 and W911NF-20-1-0163; ARO ELQ Award No. W911NF2320219; the Flatiron Institute is a division of the Simons Foundation (C.F. and S.Z.); the NSF Graduate Research Fellowship Program (L.H.K. and A.K.); the AWS Generation Q Fund at the Harvard Quantum Initiative (Y.G.); the Swiss National Science Foundation (M.L.); the Intelligence Community Postdoctoral Research Fellowship Program at Harvard administered by Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy and the Office of the Director of National Intelligence (ODNI) (A.W.Y.).
{dfigure*}
f1 [Uncaptioned image] Cooling ultracold atoms by transforming a low entropy product state to strongly-correlated states. (a), Schematic of the 2D Hubbard phase diagram. The lowest temperatures reported in [11, 6, 12, 13] (gray dashed) and the estimated temperatures in the current work (red band) are marked. (b), To reach lower entropy, we prepare a gapped band insulator (BI) in contact with a gapless metallic reservoir. Entropy flows from the BI to the reservoir [31]. After isolating the two parts, the BI is transformed into a strongly-correlated state-of-interest by dynamically changing the lattice geometry, Hubbard parameters and density, while preserving the very low entropy of the original BI. (c), The BI has a filling of two atoms per site, which appear as empty sites in a parity-projected fluorescence image (left), in contrast to the dilute reservoir. We then halve the lattice filling by doubling the number of lattice sites, thereby making the BI visible (center). The strongly-correlated state at the end of the preparation spans about $340$ sites, where empty sites correspond to coherent doublon-hole pairs indicating low entropy (right). The central state-of-interest and reservoir are shaped by optical potentials programmed with Digital Micromirror Devices (DMDs, not shown). (d), To double the number of sites, we continuously decrease the long-spacing lattice depth $V_{L}$ and increasing the short-spacing lattice depth $V_{S}$. This effectively splits doubly-occupied sites into dimers, which are then connected to form a square lattice.
{dfigure}
f2 [Uncaptioned image] Splitting a band insulator into a Heisenberg antiferromagnet. (a), Similar to the long-spacing lattice limit, the ground state at half-filling on a disconnected dimerized lattice is a product state of singlets. This state is then adiabatically connected to a long-range antiferromagnet on a square lattice by coupling nearest dimers. (b), Spin correlations $C^{zz}_{\mathbf{d}}(\mathbf{r})$ between nearest-neighbors are localized on dimers when they are weakly coupled ($\alpha=0.3$, left), and become uniform across all nearest-neighbor bonds in the square lattice limit ($\alpha=1.0$, right). (c), Spin correlations as a function of bond displacement $C^{zz}_{\mathbf{d}}$, averaged over a $r=6$ region at half-filling at $U/t=18.6(8)$. The range of the antiferromagnetic correlations starts to grow around $\alpha=0.7$. (d), Energy levels of the Heisenberg model on dimerized lattices as coupling strengths are tuned, in units of the spin exchange coupling $J=4t^{2}/U$ (see Methods). (e), The measured staggered magnetization increases at couplings $\alpha>0.7$ similar to the critical point of the quantum phase transition in the ground-state Heisenberg model [47]. Solid lines: simulation of dimerized Heisenberg model at temperature $T/J=0.5$.
{dfigure}
f3 [Uncaptioned image] Cold Hubbard antiferromagnets in the strongly-correlated regime. (a), Intermediate interaction strengths, here $U/t=8.3(2)$, lead to density fluctuations in addition to spin fluctuations at half-filling. (b), To achieve the target filling (here $n=1$), we adjust the confining potential created by DMDs to spill excess atoms. A repulsive wall isolates the cold region from the reservoir, while a volcano-shaped potential enables a spilling of excess atoms from the center. (c), Spin correlations $C^{zz}_{\mathbf{d}}$ as a function of bond displacement $\mathbf{d}$, and averaged over a radius $r=5$ region, show nearly saturated long-range antiferromagnetic correlations. The error bars denote one s.e.m. (d), Experimental data is compatible with numerically exact simulations (lines, DQMC and AFQMC) at temperatures $T/t<0.1$ and interactions $U/t=8$. (e), Low temperatures yield a sharp peak of the spin structure factor $S(\textbf{q})$ at quasi-momentum $\textbf{q}=(\pi,\pi)$. (f), We estimate temperature by computing the staggered magnetization with cutoff $d=6$ on bond distance, for both DQMC–AFQMC data (interpolated to $U/t=8.3(2)$, line) and experimental data ($1\sigma$ confidence interval shown as shaded area). Their comparison yields a temperature $T/t=0.05_{-0.05}^{0.06}$ (Methods).
{dfigure*}
f4 [Uncaptioned image] Hole-doping cold Fermi Hubbard systems. (a),(b), To coherently introduce hole dopants into the half-filled state after splitting the BI, we increase the kinetic energy and decrease the interaction strength by reducing the lattice depths out of the tight-binding regime. This prepares cold weakly interacting Fermi liquid that inherits the low initial entropy, which is adiabatically expanded to reach a given target density. Reloading the lattices and ramping up interactions create cold hole-doped Hubbard systems at $U/t=8.0(3)$ (III). Schemes of splitting the BI into a half-filled antiferromagnetic Mott insulator at strong interaction of $U/t=18.6(8)$ (I) and intermediate interaction of $U/t=8.3(2)$ (II) are also marked. (c), Doping scheme in quasimomentum space. Splitting the BI, which completely fills the Brillouin zone, into the square lattice doubles the number of sites and the size of the Brillouin zone. The population of quasimomentum states remain nearly unchanged. Expanding in real space then decreases the size of the Fermi surface and coherently introduces doping. (d) Spin correlations as a function of bond displacements $C^{zz}_{\mathbf{d}}$ measured in an ROI of $r=3$. The range and magnitude of the antiferromagnetic correlation decreases with increased doping. (e) Azimuthal average of the sign-corrected spin correlations shown in (d). At small dopings of $\delta\leq 10\%$, the antiferromagnetism still remains long-ranged over the ROI. As we increase the number of hole dopants, the strengths of spin correlations at all distances are reduced by similar amounts. (f) Short-range spin correlations $C^{zz}_{d}$ at different distances $d=1,\sqrt{2},2,\sqrt{5}$ as we dope the system. The nearest-neighbor correlations at $d=1$ show quantitative agreement between experimental data and CP-AFQMC simulations at $U/t=8$ and $T/t\in[0,0.1]$ (solid line). At longer bond distances $d>1$, experimental data show stronger correlations than CP-AFQMC simulations. The error bars denote one s.e.m.
{dfigure}
f5 [Uncaptioned image] Adiabaticity of expansion. (a) Adiabaticity of expansion from a BI to a to half-filled Hubbard system in the same lattice as a function of expansion duration $\tau$ at normalized lattice depth $\eta=0.2$. We find the ramp adiabaticity and therefore temperature improves quickly as the $\tau$ is increased from $1.76\hbar t$ (11ms), which starts to saturate at $\tau\simeq 35\hbar/t$ (2020ms). (b) Spin correlations and estimated temperatures after expansion of fermions for a fixed normalized duration measured by tunneling times $\tau\simeq 60\hbar/t$ (taking $\hbar=1$) at different lattice depths $\eta$. Expansion with $\eta\geq 0.8$ show strong heating which cannot be explained exclusively by adiabaticity. At critical lattice depth $\eta\simeq 0.6$, the interaction strength $U/t\simeq 1.5$ is similar to the interaction at which a rapid increase in transport timescales was observed in [52]. This indicates the heating may be largely dependent on a combination of interaction-induced effects and lattice heating. The error bars denote one s.e.m.
Acknowledgements
We thank Daniel Greif, Geoffrey Ji and Christie Chiu for early experimental contributions, Annabelle Bohrdt, Eugene Demler, Tilman Esslinger, Antoine Georges, Fabian Grusdt, Ehsan Khatami and Martin Zwierlein for insightful discussions, and Yuan-Yao He, Yiqi Yang and Zhou-Quan Wan for help with our calculations. We acknowledge support from the Gordon and Betty Moore Foundation, Grant No. GBMF-11521; National Science Foundation (NSF) Grants Nos. PHY-1734011, OAC-1934598 and OAC-2118310; ONR Grant No. N00014-18-1-2863; the Department of Energy, QSA Lawrence Berkeley Lab award No. DE-AC02-05CH11231; QuEra grant No. A44440; ARO/AFOSR/ONR DURIP Grants Nos. W911NF-20-1-0104 and W911NF-20-1-0163; ARO ELQ Award No. W911NF2320219; the Flatiron Institute is a division of the Simons Foundation (C.F. and S.Z.); the NSF Graduate Research Fellowship Program (L.H.K. and A.K.); the AWS Generation Q Fund at the Harvard Quantum Initiative (Y.G.); the Swiss National Science Foundation (M.L.); the Intelligence Community Postdoctoral Research Fellowship Program at Harvard administered by Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy and the Office of the Director of National Intelligence (ODNI) (A.W.Y.).
Author Contributions
M.X., L.H.K., A.K., Y.G., A.W.Y. and M.L. performed the experiment, collected and analysed data. C.F. and S.Z. performed the numerical DQMC, AFQMC and CP-AFQMC simulations, M.X. performed part of the numerical DQMC simulation in Methods, M.L. performed the Heisenberg model QMC simulations. M.G. supervised the study. All authors contributed to the interpretation of the results and production of the manuscript.
Competing Interests
M.G. is co-founder and shareholder of QuEra Computing.
{dfigure*}
exp_info [Uncaptioned image] Schematic of the experimental details used in the cooling protocol. (a). The experimental sequence of the ramps used for the optical lattice and dipole potentials to initialize a BI with ultra-low entropy. (b). The experimental sequence of the ramps used for the optical lattice and dipole potentials to split the BI into a cold antiferromagnet at half-filling in the $U/t=8.3(2)$ Hubbard model. (c). The experimental sequence of the ramps used for the optical lattice and dipole potentials to split the BI and expand into a cold doped Hubbard system at $U/t=8.3(2)$. (d). Schematic of the lattice beams. Two orthogonal laser beams $X$ and $Y$, whose phase are interferometrically stabilized, are retroreflected to create an interference lattice. A frequency-offseted laser beam $\bar{X}$, co-propagating with $X$, are also retro-reflected to create a second lattice along the $x$ direction shifted by half a site. (e). As we decrease the intensity of the $X$ beam and increase the intensity of the $\bar{X}$ beam, the interfering long-spacing lattice is ramped down and the non-interfering short-spacing lattice is ramped up. The lattice geometry changes from a long-spacing square lattice through a dimerized lattice to a short-spacing square lattice. (f). The volcano-shaped potential used in Fig. Acknowledgements. A sharp, outward sloped potential facilitates preparation of the BI but prevents adiabatic expansion of the atom cloud. (g). The flattened volcano-shaped potential used in Fig. Acknowledgements. Removing the potential barrier to form a flattened region allows adiabatic expansion. Each pixel of the DMD can be turned on or off. A continuous potential function is binarized using error diffusion algorithm [11].
{dfigure*}
exp_data [Uncaptioned image] Supplementary experimental data. (a). Correlation map and the azimuthal average computed in an ROI of $r=4$. (b). Correlation map and the azimuthal average computed in an ROI of $r=5$. (c). Using parity-projected imaging, the defect singlon density measured in the long-spacing lattice as a function of holding time. (d). Using full charge-resolved imaging, the density measured in the short-spacing lattice as a function of holding time. There is a factor of $2$ difference in the measured slope due to there are twice the number of sites. (e). Here we allow the atom cloud to expand for a constant time $\tau_{h}=120$ ms instead of constant time measured by tunneling times $\hbar/t$. The heating remain nearly the same despite the ramp time and therefore lattice heating is reduced by $\simeq 4$ times for the deepest lattice depth. The expansion time $\tau_{h}\simeq 15\hbar/t$ in the deepest lattice depth, which according to Fig. Acknowledgements should be sufficiently slow to cause negligible heating. (f). Spin correlations $C^{zz}_{d}$ shown in Fig. Acknowledgements as a function of singlon density $n_{s}$. For experimental data we correct for imaging fidelity to obtain $n_{s,\mathrm{corrected}}$ as $n_{s}$.
{dfigure*}
isolator [Uncaptioned image] Comparison between experimental methods with and without non-reciprocal attenuator at half-filling. Left to right: averages of two-dimensional singlon density, radially averaged singlon density on A/B sublattices, nearest neighbor spin correlation $C^{zz}(1)$ for intra-dimer, inter-dimer and perpendicular bonds, spin correlations $C^{zz}_{d}$ averaged over an ROI of $r=5$. The sharp edge in the spatial singlon density profile is due to the edge of the camera sensor. (a)-(d). Without the non-reciprocal attenuator, we ramp the interference time phase to $\phi=\pi/2$ to remove the tunneling dimerization. At half-filling, the resulting sublattice potential offset would not induce a density offset in the Mott insulator region but would cause offsets away from half-filling. The density offset is clearly visible in the outer reservoir region. (e)-(h). With the non-reciprocal attenuator, we could keep the interference phase at $\phi=0$. This suppresses tunneling dimerization below statistical fluctuations without inducing sublattice potential offsets. We find no difference between the densities on the two sublattices. No density modulation is detected in the reservoir.
{dfigure*}
qmc_data [Uncaptioned image] Supplementary data from numerical simulations. (a). The effects of symmetric A/B sublattice potential offset by $\pm\Delta\mu$ on spin correlations at $T/t=0.15,0.3$ using DQMC and $T/t=0$ using CP-AFQMC, with $\Delta\mu=0.0t,1.0t,2.0t$ that is symmetric $\pm\Delta\mu/2$ from 0 on A/B sublattices . We set the chemical potential $\mu=0$ at half-filling therefore there is no effect on the average density $n=1$. The densities deviation on the A/B sublattices from the average density are $\pm 0,\pm 2\%,\pm 4\%$ correspondingly. We plot the spin correlations (top) and the difference from the case with no offset (bottom). No statistically significant effect on the spin correlations is detected for $T/t\geq 0.15$ or $T/t=0,\Delta\mu\leq 1.0t$. At $T/t=0$, a potential offset of $\Delta\mu=2.0t$ seems to decrease the strength of spin correlations. Note here we use PBC and the long-range AFM order enhanced by finite size effects can be seen as saturation of spin correlation as a function of distance. (b). (Top) The doublon density $n_{d}$ as a function of density $n$ computed using CP-AFQMC at different temperatures. Below $T/t=0.15$ the variation of $n_{d}$ as a function of temperature is small compared to the statistical errors of the detected densities for all dopings where CP-AFQMC simulation is performed. (Bottom) The difference of finite temperature and ground state doublon densities $n_{d}(T)-n_{d}(0)$. (c). (Top) The doublon density $n_{d}$ as a function of temperature $T/t$ at half-filling with interaction $U/t=8$ computed using AFQMC. Below $T/t=0.2$ the variation of $n_{d}$ as a function of temperature is small compared to the statistical errors of the detected singlon densities. Below $T/t=0.15$ the $n_{d}$ is consistent with staying constant as it ground state value. (Bottom) The doublon density $n_{d}$ is sensitive to interaction strengths $U/t$ computed using DQMC for $T/t=0.1,0.15,0.2$. The saturation of $n_{d}$ as a function of temperature below $T/t=0.15$ holds for a wide range of $U/t$.
{dfigure*}
exp_qmc [Uncaptioned image] Discrepancy between CP-AFQMC results and experimental data or DQMC results. (a). Spin correlations computed using DQMC in an $8\times 8$ system and using CP-AFQMC in a $12\times 12$ system at $T/t=0.25$ and obtained from experimental data with a calibrated $U/t=8.2(2),T/t=0.264(4)$. (b). Spin correlations computed using DQMC and using CP-AFQMC in a $12\times 12$ system at $T/t=0.33$ and obtained from experimental data with a calibrated $U/t=8.2(2),T/t=0.332(7)$. The experimental data are taken using standard loading with a harmonic trap similar to Ref. [7]. We tune the atom number to reach half-filling in the center of the trap where the atoms form a large Mott insulator that can be used for thermometry. Outside the Mott insulating region, the density decrease as radial distance increases. We obtain spin correlations in the doped region by performing radial binning of the data. The singlon density $n_{s}$ is converted to density $n$ using DQMC data at $T/t=0.25$.
{dfigure*}
experiment_verify [Uncaptioned image] Supplemental experimental data to investigate potential homogeneity and verify the state to be thermalized and equilibrated. (a). Tunnelings $t_{d},t_{i},t_{p}$ as a function of $X$ lattice depth computed by fitting the band structure or constructing wannier orbitals. To balance the tunnelings within $5\%$, we need to decrease $V_{X}$ below $10^{-4}E_{R}$. DS7 is used as an example to show the density and correlation profile. (b). The spatial profile of the detected singlon density of the final two dimensional doped Hubbard system in the combined potential defined by the optical lattice and DMDs. (c). Nearest neighbor spin correlation $C^{zz}(1)$ for intra-dimer, inter-dimer and perpendicular bonds. The spin correlations are statistically consistent, confirming the systematical error caused by tunneling dimerization is small compared to statistical error. (d). Azimuthal average of the singlon density profile shown in (b) on $A/B$ sublattices. We find no difference between the densities on the two sublattices which confirms the absence of sublattice potential offset with $\phi=0$. (e). Density and spin correlations as a function of holding time before the final lattice ramp after expansion. We find no change both in the detected singlon density profile and spin correlation $C^{zz}(|\mathbf{d}|)$ for bond distances $\mathbf{d}=(1,0),(1,1),(2,0),(2,1)$. This confirms $1$ ms is sufficiently adiabatic for expansion. (f). Density and spin correlations as a function of holding time after the final lattice ramp. This confirms the system is in thermal equilibrium. (g). (Top) At full DMD potential strengths as used to prepare the BI, no statistically significant changes in density or spin correlations are measured as ramp duration is increased from $20$ ms to $50$ ms. (Bottom) At reduced DMD potential strengths, fast ramps of $20$ ms becomes no adiabatic, resulting in significant reduction in density and spin correlations. (h). Density and spin correlations as a function of holding time after the splitting ramp performed at half-filling as in Fig. Acknowledgements. The shape of density as a function of radial distances $r$ and spin correlations as a function of bond distance $d$ remain the same after holding suggests the system has thermalized and equilibrated. The reduction of the nearest-neighbor correlation may be due to lattice heating.
{dfigure*}
thermal [Uncaptioned image] Radially averaged spin correlations vs radially averaged densities. (a)-(f). Non-strict binned radially averaged spin correlations plotted against the densities of the corresponding bins, for central singlon density $n_{s}=0.744(7),0.766(7),0.800(9),0.812(9),0.855(7),0.864(8)$. Each bin contains $60$ sites, which is the same as an ROI of $r=4$.
{dfigure}
hqmc [Uncaptioned image] Numerical Quantum Monte Carlo simulation of the Heisenberg model in a dimerized square lattice. a, Bond maps of spin correlation for coupling $\alpha=0.3,0.7,1.0$ at temperature $T/J=0.5$. b, Staggered magnetization as a function of coupling and temperature. c, Sign-corrected spin correlations as a function of bond distance for experimental data at $U/t=18.6(8)$ (markers) and fit of square Heisenberg data, yielding a temperature $T/J=0.458(3)$.
{dfigure}
hf_temperature [Uncaptioned image] Systematic errors on temperature at half-filling. a, Effect of boundary conditions on DQMC simulations of the sign-corrected spin correlation function $(-1)^{d}C_{\textbf{d}}^{zz}$. b, Effect of interaction strength $U/t$ on simulated $(-1)^{d}C_{\textbf{d}}^{zz}$. c, Effect of cutoff bond distance on staggered magnetization square $(m^{z})^{2}$ and estimated temperatures. Shaded rectangles along the y- and x-axis indicate $1\sigma$ confidence intervals on measured $(m^{z})^{2}$ and estimated $T/t$, respectively. d, Effect of uncertainty on calibrated interaction strength on estimated temperatures.